OpenStudy (anonymous):

Simplify each side as much as possible, then solve the equation: 2x(x + 5) = x(x + 2) − 15

OpenStudy (anonymous):

So far I have: \[2x^2+10x=x^2=2x-15\] Do I remove the = sign?

OpenStudy (anonymous):

isolate the variable x

OpenStudy (anonymous):

+2x, not =2x

OpenStudy (anonymous):

I know the factors are: 1,2 & 3,5...I just get confused on which one should be negative.

OpenStudy (anonymous):

2x^2 + 10x = x^2 + 2x - 15 subtract x^2 from both sides so you get x^2 + 10x = 2x - 15 then subtract 2x from both sides and add 15 to both sides and u get x^2+8x+15=0

OpenStudy (anonymous):

what are multiples of positive 15 that when combined will equal positive 8?

OpenStudy (anonymous):

3 * 5 = 15 3 + 5 = 15 so you rewrite equation to look like this (x+3)(x+5)=0

OpenStudy (anonymous):

make both equal to zero and solve for x. x+3=0 subtract 3 from both sides and u get x=-3 x+5=0 subtract 5 from both sides and u get x=-5 your answer is x=-3 and x=-5

OpenStudy (anonymous):

thaaaank you, next question

OpenStudy (anonymous):

you're welcome, ok lol